41% of adults say cashews are their favorite kind of nut. You randomly sele t12adults and ask each to name his or her favorite nut. Find the probability that the number who say cashews are their favorite nut is (a) exactly three. (b) at least four, and (c) at most two. If convenient, use technology to find the probabilities. 5, (a) P(3) (b) P/x 2 4) (c) P(x s 2) Round to the nearest thousandth as needed.) (Round to the nearest thousandth as needed.) Round to the nearest thousandth as needed.)
Answer #1

Ans ) Since this is the case of binomial distribution with

p= probability of adults whose cashews is favorite kind of nut = 0.41

and q = probability of adults whose cashews is not favorite kind of nut = (1-p) = 0.59

and n= 12

we know that P(X=r) = n Cr pr(1-p)n-r

a) P(X=3) = 12C3 (0.41)3(0.59)9

P(X=3) =12C3 (0.41)3(0.59)9=0.131

b) P(X≥4) =   P(X=4)+ P(X=5)+ P(X=6)+..............+ P(X=13)

P(X≥4)=12C4 (0.41)4(0.59)9+12C5 (0.41)5(0.59)8+...............+12C3 (0.41)3(0.59)9

P(X≥4)=0.795

c)P(X≤2) =P(X=0)+ P(X=1)+ P(X=2)

=12C0 (0.41)0(0.59)12+12C1 (0.41)1(0.59)11+12C2 (0.41)2(0.59)10

P(X≤2)=0.0733